Calculate the work required to bring a charge of 5.0 static from a point 24 cm from a charge of 60 static to a point 3 cm from it.


1
Expert's answer
2021-05-13T07:19:45-0400

W=F.dr

W=Kq1*q2/r^2 dr

W=Kq1*q2(-1/r)

W=9*10^9*5*24(-1/r)

W=-1080*10^9(1/r)

When r=.24,

W=-1080*10^9(1/.24)

W=-45*10^11

When r=0.3

W=-1080*10^9*(1/0.30)

W=-36*10^11

Net work done

W=-45*10^11+36*10^11

W=-9*10^11 Nm